# Breaking the Global North Stereotype: A Global South-centric Benchmark Dataset for Auditing and Mitigating Biases in Facial Recognition Systems

### v1

@@ Introduction
Figure 4 shows the Grad-CAM activation maps for the Deepface predictions over 12 representative images from the FARFace dataset. We observe that the FRS model’s attention heatmap when classifying images as male is very different compared to when it classifies them as female. For the male classifications (the first two rows in Figure 4), the model’s focus is on a narrow vertical region stretching from the forehead to the mouth, with the most important region being the nose, as seen from the color grading on the activation map. On the other hand, when the model predicts an image as female, the region of interest varies and random pixels are highlighted in the image. The patterns are independent of the region to which the individual belongs (GN or GS). We present some representative images here, but the observations can be generalized over all images in the dataset (last column in Fig. 4).
@@
Results for few-shot learning: In Table 3, we report the results for the few-shot learning setup (results for the other combinations of adversarial variants are in the Supplementary section; see Tables 1016). These are compared against the baseline values from the pre-trained zero-shot setup on the held out test set of 480 images. In the pre-trained model, the accuracy of females is inferior to males, irrespective of the region on the ORIG as well as the RGB0.3 set. When the FRS is fine-tuned with one-shot examples, the accuracy for females improves significantly – by 30% for Global North (giving 100% accuracy) and by more than 55% for Global South in the ORIG set. The improvement in accuracy is similar for the RGB0.3 test set ($\approx$ 50%). Finally, we see that the accuracy on the RGB set improves with an increase in adversarial examples, whereas the opposite is true for the ORIG set. We see similar trends for the two-shot scenario, wherein the absolute values are lower than the one-shot scenario. Here as well, the accuracy on RGB set improves with an increasing number of RGB0.3 fine-tuning examples.
@@
Figure 5: Example activation maps from the Grad-CAM analysis of the ORIG set (held-out test set) for Deepface after two-shot fine-tuning on the ORIG set. The images are ordered as – Row 1: Males correctly predicted as male (Australia, New Zealand, Pakistan, Bangladesh), Row 2: Females incorrectly predicted as male (West Indies, Bangladesh, Pakistan, India), Row 3: Females correctly predicted as female (England, West Indies, South Africa, India) and, Row 4: Males incorrectly predicted as female (Australia, West Indies, India, England). Row 5 has the average Grad-CAM activation maps for the images of males correctly predicted as male, females correctly predicted as females and females incorrectly predicted as males; it is apparent that there is a more systematic focus on the nose when females are being correctly predicted now. No males are incorrectly identified in the held out test set.
Grad-CAM analysis: We evaluate the Grad-CAM activation maps of the held-out test set for the following setup: two-shot fine-tuning with all training examples chosen from the ORIG set. Here we observe a change in the highlighted area of the image when the model predicts the female gender – the model’s region of interest is around the nose of the person, as opposed to random areas within the image. This indicates a clear change in the model’s decision making process, specially for the true positive data points. Example activation maps are shown in Figure 5. Consider activation map in row 2, col 1 of Fig. 4 and in row 3, col 2 of Fig. 5– both belong to the same female and there is a clear change in the area of interest. This analysis generalizes over the entire held-out test set as is evident from the last row of Fig. 5 (there are no incorrectly identified males in our test set).
Results for contrastive learning: Finally, we also experiment with a novel contrastive learning setup. Recall, for the computation of the triplet loss for an anchor point (any data point in a given class) ($x_{a}$), the positive example ($x_{a}^{+}$) is an adversarial image of the same individual from RGB0.3, whereas the negative example ($x_{a}^{-}$) is an image of another person from the ORIG set of the opposite gender (other variants are discussed in the Supplementary section).
Table 6 shows the accuracy in the ORIG set for gender prediction increases significantly for females from both Global North (18.75%) and Global South (50%). The increase is more significant for images in the RGB0.3 set with an increment of 37.5% in the Global North and 60% in the Global South, thereby justifying our choice of adversarial samples in the learning setup. In compensation, the accuracy for males drops, with a max drop of $\approx 20\%$. We do not observe a large reduction in bias between the different groups here.
@@
Fairness vs. accuracy: As this is a binary classification task, we hypothesize that the drop in accuracy for males and significant improvement in accuracy for females is a result of an angular shift in the separating hyperplane, causing a small number of previously classified males to now be classified as female, but a larger number of females to be classified correctly. Thus we see that that the cost of fairness (reduction in disparity between the performance for the two genders) is a drop in accuracy for males, the majority class, which is typically a standard observation in the literature. A more in-depth mathematical study to bound this shift in accuracy, while maintaining fairness is deferred to a future study.
Takeaways: (a) The model is highly receptive to few-shot fine-tuning setups showing a significant improvement in accuracy, especially for females in the Global South, and, (b) Adversarial examples improve the accuracy as well as robustness of the fine-tuning setup.
@@
Table 7: Average country prediction accuracies for the two regions, for the fine-tuning (F) and contrastive learning (C) setup. In the fine-tuning setup, first, the model is fine-tuned on Geofaces, followed by a 2-shot fine-tuning FARFace. In the second setup, the model learned through contrastive learning first on the Geofaces dataset and followed by a second round of contrastive learning on the FARFace dataset. The two rounds of contrastive learning produce the best results.
@@
Benchmark audit of FRSs: In our benchmark audit, we see that all models have a good accuracy on the ORIG set, but commercial models are less robust to adversarial inputs; all FRSs report disparate results against Global South females (generally dark skintone). Such biases were observed previously on commercial FRSs (Buolamwini and Gebru 2018; Raji et al. 2020; Jaiswal et al. 2022). Thus we observe both temporal and emergent biases against dark-skinned women (Mehrabi et al. 2021) through our study. Our current study supports existing observations and necessitates the need for continual temporal audits of FRSs, specially for adversarial real-world inputs.
@@
Bias mitigation on Deepface: Our few-shot and contrastive learning approaches result in improved accuracy for females in both regions. Accuracy for the RGB0.3 set also improves, resulting in increased model robustness. Few shot fine-tuning performs better than contrastive learning overall, but we defer further exploration of the reasoning to a future study. Simple techniques like fine-tuning and contrastive learning improve accuracy while reducing bias, and adversarial examples improve robustness.
Limitations and scope for future work: While the current dataset has images from both Global South and Global North, the age related information and gender diversity (as the sport of cricket is male dominated) is missing. We would like to improve the dataset on these dimensions in the future. The current study focuses on classification tasks . We intend to extend it to retrieval tasks like face matching. We also plan to study the accuracy vs fairness trade-offs for our mitigation strategies and introduce better algorithms that reconciles between the two notions better.
@@
In this work, we have introduced a new large-scale face dataset curated from images of cricket players. We would like to state here that we use the gender labels already publicly available on the ESPNCricInfo website and do not attempt to modify these or any other ground truth attributes. We do not collect any private information. Our study is limited to binary gender classification, but this is primarily due to only binary gender labels – male & female being available in all face datasets (ours and in literature). We do not intend to propagate the perspective that gender is a binary label; we acknowledge that gender is a fluid spectrum, and hence, all gender labels in our study are to be interpreted as perceived gender.
Next, we note that our dataset is unbalanced, with a larger ratio of males than females. This is no surprise as it is reflective of gender imbalance observed in many professions, including sports (Singh 2023) and is a societal issue that needs deeper reflection and redressal through governmental intervention. Multiple prior open-source models have also been trained on such imbalanced datasets, as evident from their bias against females in audit studies. We do not condone this, and to prevent further misuse, we will only share a balanced subset of our dataset for any and all research purposes. Even in our bias mitigation experiments, we have only used balanced subsets (for both fine-tuning and contrastive learning). We also audit a balanced subset of images and identify that similar biases as shown here persist (Results are present in the Supplementary material).
Next, images from the Global South can be impacted with additional dimensions of disparity, such as the individuals’ socio-economic position in society and participation in only a single sport – cricket. These attributes intersect with sensitive features like skin tone and can lead to newer skews beyond Global North and South. We acknowledge this as a limitation of our dataset and recommend that future users of the dataset be mindful when performing sociotechnical analyses.
Finally, our red teaming experiment has been done to show the relative ease of misusing FRSs for controversial and morally ambiguous experiments like predicting the country of a person from their face. Ironically, the results from this experiment, too, are biased, thus showing how models trained on datasets from the Global North encode strong biases. We will not release the code or the specific subset of images used for this experiment.

### v2

@@ Introduction
Figure 4 shows the Grad-CAM activation maps for the Deepface predictions using 12 representative images from the FARFace dataset. We observe that the FRS model’s attention heatmap when classifying images as male is very different compared to when it classifies them as female. For the male classifications (the first two rows in Figure 4), the model’s focus is on a narrow vertical region stretching from the forehead to the mouth, with the most important region being the nose, as seen from the color grading on the activation map. On the other hand, when the model predicts an image as female, the region of interest varies and random pixels are highlighted in the image. The patterns are independent of the region to which the individual belongs (GN or GS). We present some representative images here, but the observations can be generalized over all images in the dataset (last column in Fig. 4).
@@
Results for few-shot learning: In Table 3, we report the results for the few-shot learning setup (more results are shared in the Supplementary section). These are compared against the baseline values from the pre-trained zero-shot setup on the held out test set of 480 images. In the pre-trained model, the accuracy of females is inferior to males, irrespective of the region on the ORIG as well as the RGB0.3 set. When the FRS is fine-tuned with one-shot examples, the accuracy for females improves significantly – by 30% for Global North (giving 100% accuracy) and by more than 55% for Global South in the ORIG set. This improvement is similar for the RGB0.3 test set ($>40\%$). Finally, we see that the accuracy for females on the RGB set improves with an increase in adversarial examples, whereas the opposite is true for the ORIG set. We see similar trends for the two-shot scenario, wherein the absolute values are lower than the one-shot scenario. Here, the accuracy on both sets improves with an increasing number of RGB0.3 fine-tuning examples.
@@
Figure 5: Example activation maps from the Grad-CAM analysis of the ORIG set (held-out test set) for Deepface after two-shot fine-tuning on the ORIG set. The images are ordered as – Row 1: Males correctly predicted as male (Australia, New Zealand, Pakistan, Bangladesh), Row 2: Females incorrectly predicted as male (West Indies, Bangladesh, Pakistan, India), Row 3: Females correctly predicted as female (England, West Indies, South Africa, India) and, Row 4: Males incorrectly predicted as female (Australia, West Indies, India, England). Row 5 has the average Grad-CAM activation maps for the images of males correctly predicted as male, females correctly predicted as females, females incorrectly predicted as males, males incorrectly predicted as females; it is apparent that there is a more systematic focus on the nose when females are being correctly predicted now.
Grad-CAM analysis: We evaluate the Grad-CAM activation maps of the held-out test set for the following setup: two-shot fine-tuning with all training examples chosen from the ORIG set. Here we observe a change in the highlighted area of the image when the model predicts the female gender – the model’s region of interest is around the nose of the person, as opposed to random areas within the image. This indicates a clear change in the model’s decision making process, specially for the true positive data points. Example activation maps are shown in Figure 5. Consider activation map in row 2, col 1 of Fig. 4 and in row 3, col 2 of Fig. 5– both belong to the same female and there is a clear change in the area of interest. This analysis generalizes over the entire held-out test set as is evident from the last row of Fig. 5.
Results for contrastive learning: Finally, we also experiment with a novel contrastive learning setup. Recall, for the computation of the triplet loss for an anchor point (any data point in a given class) ($x_{a}$), the positive example ($x_{a}^{+}$) is an adversarial image of the same individual from RGB0.3, whereas the negative example ($x_{a}^{-}$) is an image of another person from the ORIG set of the opposite gender (more results for other variants are shared in the Supplementary section). Table 6 shows the accuracy in the ORIG set for gender prediction increases significantly for females from both Global North (18.75%) and Global South (50%). The increase is more significant for images in the RGB0.3 set with an increment of 37.5% in the Global North and 60% in the Global South, thereby justifying our choice of adversarial samples in the learning setup. In compensation, the accuracy for males drops, with a max drop of $\approx 20\%$. We do not observe a large reduction in bias between the different groups here.
@@
Fairness vs. accuracy: As this is a binary classification task, we hypothesize that the drop in accuracy for males and significant improvement in accuracy for females is a result of an angular shift in the separating hyperplane, causing a small number of previously classified males to now be classified as female, but a larger number of females to be classified correctly. Thus we see that the cost of fairness (reduction in disparity between the performance for the two genders) is a drop in accuracy for males, the majority class, which is typically a standard observation in the literature. A more in-depth mathematical study to bound this shift in accuracy, while maintaining fairness is deferred to a future study.
Takeaways: (a) The model is highly receptive to few-shot fine-tuning setups showing a significant improvement in accuracy, especially for females in the Global South, and, (b) Adversarial examples improve the accuracy as well as the robustness of the fine-tuning setup.
@@
Table 7: Average country prediction accuracies for the two regions, for the fine-tuning (F) and contrastive learning (C) setup. In the fine-tuning setup, first, the model is fine-tuned on Geofaces, followed by a 2-shot fine-tuning on FARFace. In the second setup, the model is trained using contrastive learning first on the Geofaces dataset and followed by a second round of contrastive learning on the FARFace dataset. The two rounds of contrastive learning produce the best results.
@@
Benchmark audit of FRSs: In our benchmark audit, we see that all models have a good accuracy on the ORIG set, but commercial models are less robust to adversarial inputs; all FRSs report disparate results against Global South females (generally dark skin tone). Such biases were observed previously on commercial FRSs (Buolamwini and Gebru 2018; Raji et al. 2020; Jaiswal et al. 2022). Thus we observe both temporal and emergent biases against dark-skinned women (Mehrabi et al. 2021) through our study. Our current study supports existing observations and necessitates the need for continual temporal audits of FRSs, especially for adversarial real-world inputs.
@@
Bias mitigation on Deepface: Our few-shot and contrastive learning approaches result in improved accuracy for females in both regions. Accuracy for the RGB0.3 set also improves, resulting in increased model robustness. Few-shot fine-tuning performs better than contrastive learning overall, but we defer further exploration of the reasoning to a future study. Simple techniques like fine-tuning and contrastive learning improve accuracy while reducing bias, and adversarial examples improve robustness.
Limitations and scope for future work: While the current dataset has images from both Global South and Global North, the age related information and gender diversity (as the sport of cricket is male dominated) is missing. We would like to improve the dataset on these dimensions in the future. The current study focuses on classification tasks. We intend to extend it to retrieval tasks like face verification. We also plan to study the accuracy vs fairness trade-offs for our mitigation strategies and introduce better algorithms that reconciles between the two notions better.
Acknowledgements
This research is supported partly by a European Research Council (ERC) Advanced Grant for the project “Foundations for Fair Social Computing”, funded under the European Union’s Horizon 2020 Framework Programme (grant agreement no. 789373), and by a grant from the Max Planck Society through a Max Planck Partner Group at IIT Kharagpur. S. Jaiswal is supported by the Prime Minister’s Research Fellowship, Govt. of India.
@@ Acknowledgements
In this work, we have introduced a new large-scale face dataset, curated from images of cricket players. We would like to state here that we use the gender labels already publicly available on the ESPNCricInfo website and do not attempt to modify these or any other ground truth attributes. We do not collect any private information. Our study is limited to binary gender classification, but this is primarily due to only binary gender labels – male & female being available in all face datasets (ours and in literature). We do not intend to propagate the perspective that gender is a binary label; we acknowledge that gender is a fluid spectrum, and hence all gender labels in our study are to be interpreted as perceived gender.
Next, we note that our dataset is unbalanced, with a larger ratio of males than females. This is no surprise as it is reflective of gender imbalance observed in many professions, including sports (Singh 2023) and is a societal issue that needs deeper reflection and redressal through governmental intervention. Multiple prior open-source models have also been trained on such imbalanced datasets as is evident from their bias against females in audit studies. We do not condone this and, as such, to prevent further misuse, will only share a balanced subset of our dataset for any and all research purposes. We also audit on a balanced subset of images and identify that similar biases as shown here persist (Results are in an extended version of the paper66footnotemark: 6). Even in our bias mitigation experiments, we have only used balanced subsets (for both fine-tuning and contrastive learning).
Images from the Global South can include additional sensitive features, for example, caste or religion in South Asia, which are correlated to one’s socio-economic position in these societies. These attributes intersect with sensitive features like skin tone and can lead to newer skews beyond Global North and South. We acknowledge this as a limitation of our dataset and recommend future users of the dataset to be mindful when performing sociotechnical analyses.
Finally, our red teaming experiment has been done to show the relative ease of misusing FRSs for controversial and morally ambiguous experiments like predicting the country of a person from their face. Ironically, the results from this experiment too are biased, thus showing how models trained on datasets from the Global North encode strong biases. We will not be releasing either the code or the specific subset of images used for this experiment.